LBTO Technial/D/Service Night

OSA: D.G.Huerta

ISp: J.Morris

Manager: Michael Wagner

ISA: B.Rothberg, D.Thompson (Tucson)

=======================================

UT 12:52 - Enclosure open

UT 01:03 - Collimating - seeing <1" in guider

UT 01:04 - M015StoneO MOS test

UT 01:08 - Acq failed - no guidestar found

UT 01:14 - on guidestar but LUCI-2 doesn't seem to be doing anything - Running box vanished in GUI

UT 01:19 - Nothing happening in LUCI-2, so putting mask in FPU - reinitializing all

UT 01:21 - resent preset after collimation

UT 01:25 - reset and re-sent preset - took quick image with GEIRS - all is ok there

UT 01:27 - LUCI-2 is doing nothing - Initialized all again

UT 01:40 - restarted LUCI-2 scheduler

UT 02:09 - did first MOS acquisition

UT 02:10 - did test spectra (MOS test 0001-0007)

======================================================

UT 02:12 - Moving to INAF Beltran A1 G35.20-0.74N Ks-imaging (conditions meet criteria)

UT 02:14 - taking science first exposure luci2.20151101.0008-0027

UT 02:18 - first pair subtraction images look elongated - airmass 1.517

UT 02:25 - Snoopy dance

UT 02:26 - ambient outside temperature suddenly increasing - this is affecting seeing

UT 02:54 - A1 G35.20-0.74N H2 filter - first exposure 0028

UT 03:05 - median guider 0.7"

UT 03:22 - median guider 0.8"

UT 03:49 - 16/30 completed H2 - last exposure 0043

UT 03:51 - Coordinates for standard FS148 in script were wrong. They were for SJ875C so I could not get a standard.

UT 03:55 - Will try to get another standard soon

==========================================

UT 03:55 - Moving to Dual AGN target SDSSJ2356-1016

UT 03:58 - I loaded AGN xml file after pausing and then aborting imaging script. Scheduler got lost and decided to continue H2 imaging of Dual AGN spectroscopy script (0044)

UT 04:00 - Restarting Scheduler

UT 04:10 - had to redo script for Dual AGN with correct PA (0045-0050 bad)

UT 04:19 - Re-running SDSSJ2356-1016

UT 04:22 - accidentally clicking subtract image in Longslit tab on RTD2 froze observer gui

UT 04:30 - had to edit PA to 173 from 146 - still not right

============================================

UT 04:37 - Moving to MWagner Assassin field JHK imaging (0057-0071)

UT 04:56 - 0.42" seeing GUIDER

UT 05:04 - Assassin field complete

====================================

UT 05:05 - Moving to standard FS29 for INAF Beltran program

UT 05:08 - Scriptor still kept old coordinates - made V2 with Ks and H2 filters for FS 29

UT 05:13 - Had to abort - GS not in field - doing a pointing check

UT 05:15 - bright crud at bottom of guider is catching the GS even though star is 13.43 -> star is there, guider is just going to bright crud at bottom of ccd

UT 05:18 - Sending preset again

UT 05:22 - changed guidestar and trying again

UT 05:25 - FS29 standard airmass 1.575 Ks,H2 (0072-0081)

UT 05:38 - FS29 complete - airmass 1.660

==================================

UT 05:39 - Moving to SDSSJ2356-1016 again (0082-0113)

UT 05:47 - Going to do two acquisitions and two runs on script instead of wasting time putting both targets in slit; PA=50 - this shows targets next to each other - SDSS at PA=0 looks like its actually PA 90

UT 05:50 - starting spectroscopy

UT 06:14 - re-acquiring in same field - will center up on 2nd galaxy

UT 06:16 - Error in commit - instrument manager setup

UT 06:16 - sending preset again

UT 06:21 - seeing about 0.6" in guider

UT 06:24 - Real Time Display lost track of which is the longslit

UT 06:26 - Center in the slit seems off by 3 or more pix- edge detection issue?

UT 06:28 - Centered and starting spectra again

UT 06:30 - Error during commit - error commiting to instrument manager again

UT 06:32 - Aborting and restarting scheduler as LUCI-2 doesn't know what to do

UT 06:33 - Resending preset after restarting scheduler

UT 06:41 - Longslit acquisition center slit position is pulling towards the left again, wants to center at 1016 instead of ~1021 which is actual center of slit

UT 06:46 - observer gui froze again

UT 06:49 - Had to add detxy +0.5" to center in slit

UT 06:51 - Resending acquisition - centering is completely off, object is at Y=972

UT 06:59 - centering longslit again put it at 1017 - going with it to see if it actually works ok

UT 07:01 - looks like it did put it off center by 2-3 pixels - doing 0.25" (2pix) +detxy move

UT 07:04 - Bang ON! correction i made was right - starting integration

UT 07:22 - RTD2 slow on subtracting two images ~ 30 sec to complete

UT 07:28 - Dual AGN complete - moving to HD1154 Telluric

UT 07:32 - LS alignment again - center slit position is centering at 1019 instead of 1021

UT 07:34 - airmass 1.92 - my selection is correct, not 1019

UT 07:35 - spectra started

UT 07:38 - DONE
====================================================

UT 07:43 - D.Miller testing transforms now - configuring for N30_oldfieldstrop N375, BrGamma

UT 07:54 - Stone_D_new_Oct27.list

UT 08:08 - put in old N30fieldstop - my bad - should be pointing mask - switching now

01140-0118 bad don't use

UT 08:44 - D.Miller working on elongated images

UT 08:52 - luci2.20151101.121 - first image with star - "center" position

UT 08:53 - 0122 - "blank"

UT 09:10 - stopped and exited

UT 09:12 - restarted transform_collect,start=10 (GS24) 0132

UT 09:12 - 0137 GS 4

UT 09:22 - 0138 GS12

UT 09:27 - 0139 GS23 PA15

UT 09:29 - 0140 GS13 PA15

UT 09:30 - 0141 GS22 PA15

UT 09:31 - 0142 GS17 PA15

UT 09:34 - 0143 GS5 PA15

UT 09:34 - 0144 GS11 PA15

UT 09:36 - 0145 GS29 PA15

UT 09:38 - 0146 GS24 PA15

UT 09:39 - 0147 GS30 PA15

UT 09:40 - 0148 GS4 PA30

UT 09:43 - 0149 GS12 PA30

UT 09:45 - 0150 GS23 PA30

UT 09:46 - 0151 GS13 PA30

UT 09:49 - 0152 GS22 PA30

UT 09:50 - 0153 GS5 PA30

UT 09:51 - 0154 GS11 PA30

UT 09:54 - "saving" taking over a minute

UT 09:56 - GEIRS is frozen

UT 09:57 - Correction - connection to mountain is frozen - lost polycom, lost connection to OBS2

UT 09:5X - 0155 GS29 PA30 -unknown if this got done

UT 10:58 - still no word on mountain

UT 12:15 - spoke on phone with summit, things still down

UT 12:25 - Observing concluded - T. Sargent and D. Cox working with DG. Huerta to restore systems on the mountain. B. Rothberg leaving remote ops, will stay in contact for a bit longer to update situation if needed.

=====================================

-- %USERSIG{BarryRothberg - 2015-11-01}%

Comments

 
Topic revision: r3 - 01 Nov 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback